# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 8.3.1 MRL  for  IPA
[…]
@@
Summary: The objective of MRL methods for IPA is to accurately and completely retrieve reference for each query, which requires robustly measure the similarity between query and reference against shifted viewpoints, changed appearance, dynamic occlusions, and so on. The utilization of pre-trained feature from computer vision community keetha2023anyloc , geometric verification based on local features yue2 , robust representation combined local and global features fild2 , temporal enhancement Bampis2018IJRR may be helpful ways to improve performance for this field. But since VPR and I2P-VPR methods directly use retrieved reference pose as current pose, their localization accuracy is inherently limited, which will be experimental concluded in later paragraphs. So they are generally used as coarse or initial pose approximation in hierarchical localization framework hloc2 .
@@
Additionally, VPR-Bench [^5] https://github.com/MubarizZaffar/VPR-Bench vprbench is an open-sourced VPR evaluation framework with quantifiable viewpoint and illumination invariance, enabling fast and full quantified evaluation for VPR methods. In I2P-VPR field, such public evaluation benchmark is not proposed yet.
Figure 17: Recall over top-K candidates performance of I2P-Rec Li2023I2PRecRI on KITTI dataset KITTI .
8.3.2 MRL  for  FPE
@@
Several MRL methods tend to estimate fine pose. They utilize various kind of scene map and achieve different level of localization accuracy. We first provide performance of RPR, VL-MRL, I2P-MRL, HD-MRL, and NN-MRL respectively. Then, most kind of VL-MRL methods are evaluated on two typical dataset so that their localization accuracy can be fairly compared.
@@
RPR: Towards evaluation of RPR methods, a benchmark called mapfree-reloc-benchmark has been recently proposed  [^6] https://research.nianticlabs.com/mapfree-reloc-benchmark/leaderboard, we suggest readers to refer to the public leader board for detailed information. Up to now, the SOTA solution of RPR is solving pose by decomposing essential matrix between matched features across query image and reference image and then recover absolute scale by a SOTA monocular depth estimation method, DPT DPT . This solution can achieve a localization performance with 1.23m translation error and 11.1° rotation error.
@@
SIFT sift 5K + NN - 20.4 / 27.6 / 37.8
SIFT + CAPS CAPS + NN - 77.6 / 86.7 / 99.0
SuperPoint superpoint + NN - 73.5 / 79.6 / 88.8
SuperPoint + S2DNet S2DNet - 74.5 / 84.7 / 100.0
SuperPoint + CAPS CAPS + NN - 82.7 / 87.8 / 100.0
D2-Net (SS) d2net 1K + NN - 64.3 / 78.6 / 91.8
D2-Net (SS) d2net 2K + NN - 74.5 / 85.7 / 96.9
D2-Net (SS) d2net 5K + NN - 77.6 / 92.9 / 100.0
D2-Net (MS) d2net 1K + NN - 53.1 / 74.5 / 86.7
D2-Net (MS) d2net 2K + NN - 61.2 / 81.6 / 94.9
R2D2 r2d2 1K + NN - 55.1 / 70.4 / 77.6
R2D2 r2d2 2K + NN - 63.6 / 78.6 / 87.8
R2D2 r2d2 5K + NN - 69.4 / 84.7 / 98.0
ASLFeat aslfeat + NN - 77.6 / 89.8 / 100.0
SEKD (SS) sekd 1K + NN - 30.6 / 33.7 / 38.8
SEKD (SS) sekd 2K + NN - 42.9 / 51.0 / 57.1
SEKD (SS) sekd 5K + NN - 56.1 / 69.4 / 79.6
SEKD (MS) sekd 1K + NN - 35.7 / 42.9 / 50.0
SEKD (MS) sekd 2K + NN - 50.0 / 63.3 / 70.4
SEKD (MS) sekd 5K + NN - 67.3 / 74.5 / 83.7
DISK disk 1K + NN - 60.2 / 72.4 / 81.6
DISK disk 2K + NN - 70.4 / 82.7 / 94.9
DISK disk 5K + NN - 80.6 / 86.7 / 99.0
GLFeat glfeat 5K + NN - 79.6 / 86.7 / 100.0
ALIKE-L alike 1K + NN - 59.2 / 73.5 / 83.7
ALIKE-L alike 2K + NN - 74.5 / 87.8 / 98.0
ALIKED-N(32) aliked 1K + NN - 77.6 / 88.8 / 100.0
ALIKED-N(32) aliked 2K + NN - 76.5 / 87.8 / 100.0
DualRC-Net DRCNet - 79.6 / 88.8 / 100.0
SparseNCNet SparseNCNet - 76.5 / 84.7 / 98.0
Patch2Pix patch2pix - 79.6 / 87.8 / 100.0
Full Localization with HLoc hloc2
SuperPoint superpoint + NN 85.4 / 93.3 / 97.2 75.5 / 86.7 / 92.9
SuperPoint + CAPS CAPS + NN 86.3 / 93.0 / 95.9 83.7 / 90.8 / 96.9
SuperPoint + SuperGlue superglue 89.6 / 95.4 / 98.8 86.7 / 93.9 / 100.0
Patch2Pix patch2pix 84.6 / 92.1 / 96.5 82.7 / 92.9 / 99.0
Patch2Pix (w.CAPS) patch2pix 86.7 / 93.7 / 96.7 85.7 / 92.9 / 99.0
Patch2Pix (w.SuperGlue) patch2pix 89.2 / 95.5 / 98.5 87.8 / 94.9 / 100.0
@@
    Some results are cited from Patch2Pix patch2pix and we do not know the limitation of number of extracted features or matched features. The results with limited number of features are denotes like “SIFT 5K” and cited from ALIKED aliked and GLFeat glfeat .
Table 4: Localization success ratio of some VL-MRL methods on Aachen Day-Night Benchmark (v1.0) benchmark6dof .
Table 5: Localization success ratio of some VL-MRL methods on InLoc Benchmark  InLoc .
@@
SuperPoint superpoint + NN 40.4 / 58.1 / 69.7 42.0 / 58.8 / 69.5
D2Net d2net + NN 38.4 / 56.1 / 71.2 37.4 / 55.0 / 64.9
R2D2 r2d2 + NN 36.4 / 57.6 / 74.2 45.0 / 60.3 / 67.9
SuperPoint + SuperGlue superglue 49.0 / 68.7 / 80.8 53.4 / 77.1 / 82.4
SuperPoint + CAPS CAPS + NN 40.9 / 60.6 / 72.7 43.5 / 58.8 / 68.7
SIFT + CAPS CAPS + NN 38.4 / 56.6 / 70.7 35.1 / 48.9 / 58.8
SparseNCNet SparseNCNet 41.9 / 62.1 / 72.7 35.1 / 48.1 / 55.0
Patch2Pix patch2pix 44.4 / 66.7 / 78.3 49.6 / 64.9 / 72.5
Patch2Pix (w.SuperPoint+CAPS) patch2pix 42.4 / 62.6 / 76.3 43.5 / 61.1 / 71.0
Patch2Pix (w.SuperGlue) patch2pix 50.0 / 68.2 / 81.8 57.3 / 77.9 / 80.2
Table 6: 6 DoF localization errors of some I2P-MRL methods on KITTI Odometry Dataset KITTI
@@
CMRNet cmrnet 1.09 0.26 1.39 0.51
HyperMap hypermap n.a. n.a. 1.42 0.48
POET miao2023poses 0.91 0.25 0.79 0.20
I2PNet I2PNet 0.74 0.08 0.67 0.07
@@
VL-MRL: As for VL-MRL methods, as shown in Tab. 4 and Tab. 5, many recent SOTA method in local feature extraction superpoint ; disk ; CAPS ; aslfeat and matching superglue ; DRCNet ; SparseNCNet can be readily integrated into HLoc framework as alternates to originally used HFNet hloc2 and NN matching. We report a widely used evaluation metric here, which is the ratio of successfully localized queries within three defined tolerances of localization errors. If there were enough correctly matched local features between query image and visual landmark map, HLoc can achieve an excellent localization performance. From Tab. 4, it can be seen that extracting more local features contribute to localization, but MS detection is not constantly positive for localization d2net . The same key point detectors with more advanced descriptors will generate more accurate matches so they can achieve better performance S2DNet ; CAPS .
From Tab. 4 and 5, we can also see that accurate localization in indoor scene InLoc is much more difficult than localization in urban scene benchmark6dof due to lack of textures. In current framework of VL-MRL methods that local features are widely used as scene representation, textures of scene still be necessary.
I2P-MRL: Tab. 6 shows some cross-modal MRL methods on KITTI 00 sequence KITTI . The learning-based I2P-MRL methods cmrnet ; cmrnet++ ; hypermap ; miao2023poses ; I2PNet ; I2DLoc seem to outperform geometry-based one mono_i2p_vo with a lot of margin. The SOTA I2P-MRL methods can achieve an impressive localization performance with 0.08 m in translation error and 0.74°in rotation error I2PNet . However, these I2P-MRL methods are trained and tested on human-defined data that the displacement between query image and reference point cloud is randomly determined within a defined threshold and it is often limited. So, their performance in practical applications is still waiting for evaluation.
Table 7: 3 DoF Localization errors of some HD-MRL methods.
@@
Choi et al. Choi2019Access 16km highways 0.12 0.18 n.a.
@@
Xiao et al. Xiao2018TISC KITTI 04 KITTI 0.345
Liao et al. Liao2019CoarseToFineVL KAIST 26 KAIST 0.64 0.90
Virtual KITTI gaidon2016virtual 0.29 0.32
MLVHM MLVHM Campus 0.21
@@
ROADMap RoadMap 22km Urban 0.04 0.04 0.12
Wang et al. Wang2021ICRA Urban canyons 0.46 0.13 0.11
@@
KAIST 23 KAIST 0.32 0.05 0.61
KAIST 26 KAIST 0.16 0.05 0.48
@@
Zhang et al. Zhang2023TIV Highway 0.17 0.04 n.a.
BEV-Locator bevlocator nuScenes nuscenes2019 0.18 0.08 0.51
@@
EgoVM he2023egovm nuScenes nuscenes2019 0.11 0.03 0.09
@@
HD-MRL: In Tab. 7, we collect the results of some HD-MRL methods from their publications. Since the public HD map data is very few in existing dataset, researchers often collect and build their own dataset to evaluate the proposed methods. For fair and full comparison, we also list the evaluated scene/dataset for each method. It can be seen that the tightly coupling of VO and HD-MRL TM3Loc improves the accuracy in long-term localization. The learning-based methods effectively solve the data association problem between 2D images and HD map elements, thus outperforming traditional methods, and they can localize without the requirements of online VO threads. BEV-Locator he2023egovm estimates pose by regression whereas EgoVM he2023egovm relies on exhaustively searching pose candidates, the latter solution achieves better performance. It should be noticed that HD map is a very compact format of map representation. In MLVHM MLVHM , the HD maps stored in ASCII format only requires about 50 KB storage space per kilometer, compared to about 600 MB per kilometer of the raw point cloud map. It is not the only case that the map size of EgoVM is 0.35MB/km compared to 5.92MB/km of DA4AD DA4AD . Therefore, the compactness and accuracy of HD map makes it a popular choice for localization in AD.
Table 8: Localization results on Cambridge Landmarks dataset posenet .
@@
VLAD tokyo247 2.80,5.7° 4.01,7.1° 1.11,7.6° 2.31,8.0°
VLAD+Inter Sattler2019UnderstandingTL 1.48,4.5° 2.68,4.6° 0.90,4.3° 1.62,6.1°
@@
EssNet EssNet 0.76,1.9° 1.39,2.8° 0.84,4.3° 1.32,4.7°
NC-EssNet EssNet 0.61,1.6° 0.95,2.7° 0.70,3.4° 1.12,3.6°
RelocGNNDSAC2 0.48,1.0° 1.14,2.5° 0.48,2.5° 1.52,3.2°
@@
PoseNet posenet 1.92,5.40° 2.31,5.38° 1.46,8.08° 2.65,8.48°
BayesianPN BayesianPN 1.74,4.06° 2.57,5.14 ° 1.25,7.54° 2.11,8.38°
LSTM-PN LSTMPN 0.99,3.65° 1.51,4.29° 1.18,7.44° 1.52,6.68°
SVS-Pose SVS-Pose 1.06,2.81° 1.50,4.03° 0.63,5.73° 2.11,8.11°
GPoseNet Cai2018AHP 1.61,2.29° 2.62,3.89° 1.14,5.73° 2.93,6.46°
PoseNetLearn PoseNet2 0.99,1.06° 2.17,2.94 ° 1.05,3.97° 1.49,3.43°
GeoPoseNet PoseNet2 0.88,1.04° 3.20,3.29° 0.88,3.78° 1.57,3.32°
MapNet mapnet 1.07,1.89° 1.94,3.91° 1.49,4.22° 2.00,4.53°
IRPNet Shavit2020DoWR 1.18,2.19° 1.87,3.38° 0.72,3.47° 1.87,4.94°
@@
MSPN MSPN 1.73,3.65° 2.55,4.05° 2.92,7.49° 2.67/6.18°
MS-TransMS-Trans1 0.83,1.47° 1.81,2.39° 0.86,3.07° 1.62,3.99°
c2f-MS-Trans MS-Transformer2 0.70,2.69° 1.48,2.94 ° 0.59,2.88° 1.14,3.88°
@@
DSAC DSAC 0.30,0.5° 0.33,0.6° 0.09,0.4° 0.55,1.6°
DSAC++ Less_is_More 0.18,0.3° 0.20,0.3° 0.06,0.3° 0.13,0.4°
$*{}^{*}$ DSAC2 0.18,0.3° 0.21,0.4° 0.05,0.3° 0.15,0.5°
SANet SANet 0.32/0.5° 0.32/0.5° 0.10/0.5° 0.16/0.6°
@@
DFNet DFNet n.a. 0.46,0.87° 0.16,0.59° 0.50,1.49°
LENS Moreau2021LENSLE 0.33,0.5° 0.44,0.9° 0.27,1.6° 0.53,1.6°
@@
NeRF-Loc liu2023nerfloc 0.11,0.2° 0.18,0.4° 0.04,0.2° 0.07,0.2°
Table 9: Localization results on 7Scenes dataset posenet .

### v3

@@ VIII-B2 MRL  for  FPE
VIII-C Performance of SOTAs
VIII-C1 MRL  for  CPA
MRL methods for CPA, including VPR and I2P-VPR, tend to retrieve reference data given query image, so they are measured and compared based on the precision and recall metrics of their retrieved results.
VPR: First, we give some performance of VPR methods in public datasets, evaluated with different metrics, e.g., R@K (in Tab. III) and maxP@R=1 (in Tab. IV). It shows in Tab. III that the differentiable VLAD pooling layer could effectively aggregate dense local descriptors into global features, thus making NetVLAD [19] an excellent solution in retrieval-based GF-VPR tasks. Then, benefit from the global fusion of features, MixVPR [66] achieves improved performance, espeically in challenging Nordland dataset [289] with varying weather and seasons. More recently, pre-trained features from visual foundation model (such as DINOv2 [67]) are concluded to be an excellent alternate for GF-VPR, and enable AnyLoc [42] achieves impressive retrieval performance in very different scene. Although GF-VPR methods can fully retrieve reference images, their retrieval accuracy is slightly lower than SOTA LF-VPR methods as shown in Tab. IV. Adding verification with local features [88] or jointly utilizing local feature and global feature [74] can achieve pleased accuracy while retrieve most of reference images for each query image.
I2P-VPR: As for cross-modal I2P-VPR between query image and reference point cloud clip, a SOTA performance is achieved by I2P-Rec [200] as shown in Fig. 16. With the assistance of mature monocular depth estimation, I2P-Rec [200] transform query image and point cloud clip from scene map to BEV images, simplifying the cross-modal retrieval problem and leading to pleased performance.
Figure 16: Recall over top-K candidates of I2P-Rec [200] on KITTI [288].
@@ VIII-C1 MRL  for  CPA
Summary: The objective of MRL methods for CPA is to accurately and completely retrieve reference for each query, which requires robustly measure the similarity between query and reference against shifted viewpoints, changed appearance, dynamic occlusions, and so on. The utilization of pre-trained feature from computer vision community [42], geometric verification based on local features [88], robust representation combined local and global features [74], temporal enhancement [299], and transferring into unified representation space [200] may be helpful ways to improve performance for this field. But VPR and I2P-VPR methods directly use retrieved reference pose as current pose, their localization accuracy is inherently limited, which will be experimental concluded in later paragraphs. So they are generally used as coarse pose guess in hierarchical localization framework.
@@
Additionally, VPR-Bench [^5] https://github.com/MubarizZaffar/VPR-Bench [15] is an open-sourced VPR evaluation framework with quantifiable viewpoint and illumination invariance, enabling fast and full quantified evaluation for VPR methods. In I2P-VPR field, such public evaluation benchmark is not proposed yet.
VIII-C2 MRL  for  FPE
@@
Several MRL methods tend to estimate fine pose. They utilize various kind of scene map and achieve different level of localization accuracy. We first provide performance of RPR, VL-MRL, I2P-MRL, HD-MRL, and NN-MRL respectively. Then, these methods are evaluated on two datasets [268, 27] so that they can be fairly compared.
@@
RPR: Towards evaluation of RPR methods, a benchmark called mapfree-reloc-benchmark has been recently proposed  [^6] https://research.nianticlabs.com/mapfree-reloc-benchmark/leaderboard, we suggest readers to refer to the public leader board for detailed information. Up to now, the SOTA solution of RPR is solving pose by decomposing essential matrix between matched features across query image and reference image and then recover scale by a monocular depth estimation method [109]. It can achieve a localization performance with 1.23m translation error and 11.1° rotation error.
@@
SIFT [56] 5K - 20.4 / 27.6 / 37.8
SIFT + CAPS [300] - 77.6 / 86.7 / 99.0
SuperPoint [57] - 73.5 / 79.6 / 88.8
D2-Net (SS) [130] 1K - 64.3 / 78.6 / 91.8
D2-Net (SS) [130] 5K - 77.6 / 92.9 / 100.0
D2-Net (MS) [130] 1K - 53.1 / 74.5 / 86.7
R2D2 [97] 1K - 55.1 / 70.4 / 77.6
R2D2 [97] 5K - 69.4 / 84.7 / 98.0
ASLFeat [126] - 77.6 / 89.8 / 100.0
SEKD (SS) [129] 1K - 30.6 / 33.7 / 38.8
SEKD (SS) [129] 5K - 56.1 / 69.4 / 79.6
SEKD (MS) [129] 1K - 35.7 / 42.9 / 50.0
SEKD (MS) [129] 5K - 67.3 / 74.5 / 83.7
DISK [94] 1K - 60.2 / 72.4 / 81.6
DISK [94] 5K - 80.6 / 86.7 / 99.0
GLFeat [136] 5K - 79.6 / 86.7 / 100.0
ALIKE-L [95] 1K - 59.2 / 73.5 / 83.7
ALIKED-N(32) [96] 1K - 77.6 / 88.8 / 100.0
DualRC-Net [159] - 79.6 / 88.8 / 100.0
SparseNCNet [158] - 76.5 / 84.7 / 98.0
Patch2Pix [301] - 79.6 / 87.8 / 100.0
Full Localization with HLoc [24]
SuperPoint [57] 85.4 / 93.3 / 97.2 75.5 / 86.7 / 92.9
SuperPoint +   SuperGlue [98] 89.6 / 95.4 / 98.8 86.7 / 93.9 / 100.0
Patch2Pix [301] 84.6 / 92.1 / 96.5 82.7 / 92.9 / 99.0
Patch2Pix   (w.CAPS) [301] 86.7 / 93.7 / 96.7 85.7 / 92.9 / 99.0
Patch2Pix   (w.SuperGlue) [301] 89.2 / 95.5 / 98.5 87.8 / 94.9 / 100.0
@@
    Some results are cited from [301] and we do not know the limitation of number of features. The results with limited number of features are denotes like “SIFT 5K” and cited from [96, 136].
TABLE V: Localization success ratio of some VL-MRL methods on Aachen Day-Night Benchmark (v1.0) [33].
VL-MRL: As for VL-MRL methods, as shown in Tab. V and Tab. VI, we report the ratio of successfully localized queries within three pre-defined tolerances of localization errors.. Many recent SOTA method in local feature extraction [57, 94, 300, 126] and matching [98, 159, 158] can be readily integrated into HLoc framework as alternates to originally used HFNet [24] and NN matching. If there were enough correctly matched local features between query image and visual landmark map, HLoc can achieve an excellent localization performance. From Tab. V, it can be seen that extracting more local features contribute to localization, but MS detection is not constantly positive for localization [130]. The same local features with better feature matching methods will generate more accurate matches so they can achieve better performance [98]. From Tab. V and VI, we can also see that accurate localization in indoor scene [35] is much more difficult than localization in urban scene [33] due to repeated textures or textureless regions. In current framework of VL-MRL methods that local features are widely used as scene representation, textures of scene still be necessary.
TABLE VI: Localization success ratio of some VL-MRL methods on InLoc Benchmark  [35], cited from [301].
@@
SuperPoint [57] 40.4 / 58.1 / 69.7 42.0 / 58.8 / 69.5
D2Net [130] 38.4 / 56.1 / 71.2 37.4 / 55.0 / 64.9
R2D2 [97] 36.4 / 57.6 / 74.2 45.0 / 60.3 / 67.9
SuperPoint   + SuperGlue [98] 49.0 / 68.7 / 80.8 53.4 / 77.1 / 82.4
SparseNCNet [158] 41.9 / 62.1 / 72.7 35.1 / 48.1 / 55.0
Patch2Pix [301] 44.4 / 66.7 / 78.3 49.6 / 64.9 / 72.5
Patch2Pix   (w.SuperGlue) [301] 50.0 / 68.2 / 81.8 57.3 / 77.9 / 80.2
TABLE VII: 6 DoF localization errors of some I2P-MRL methods on KITTI Odometry Dataset [288]
@@
CMRNet [18] 1.09 0.26 1.39 0.51
HyperMap [201] n.a. n.a. 1.42 0.48
POET [202] 0.91 0.25 0.79 0.20
I2PNet [211] 0.74 0.08 0.67 0.07
@@
I2P-MRL: Tab. VII shows some cross-modal MRL methods on KITTI 00 sequence [288]. The learning-based I2P-MRL methods [18, 205, 201, 202, 211, 209] seem to outperform geometry-based one [17] with a lot of margins. The SOTA I2P-MRL methods can achieve an impressive localization performance with 0.08 m in translation error and 0.74°in rotation error [211]. However, these I2P-MRL methods are trained and tested on human-defined data that the displacement between query image and reference point cloud is randomly determined within a defined threshold and it is often limited. So, their performance in practical applications is still waiting for evaluation.
TABLE VIII: 3 DoF Localization errors of some HD-MRL methods.
@@
Choi et al. [302] 16km highways 0.12 0.18 n.a.
@@
Xiao et al. [225] KITTI 04 [288] 0.345
Liao et al. [226] KAIST 26 [287] 0.64 0.90
Virtual KITTI [303] 0.29 0.32
MLVHM [9] Campus 0.21
@@
ROADMap [11] 22km Urban 0.04 0.04 0.12
Wang et al. [227] Urban canyons 0.46 0.13 0.11
@@
KAIST 23 [287] 0.32 0.05 0.61
KAIST 26 [287] 0.16 0.05 0.48
@@
Zhang et al. [304] Highway 0.17 0.04 n.a.
BEV-Locator [233] nuScenes [305] 0.18 0.08 0.51
@@
EgoVM [234] nuScenes [305] 0.11 0.03 0.09
@@
HD-MRL: In Tab. VIII, we collect the results of some HD-MRL methods from their publications. Since the public HD map data is very few in existing dataset, researchers often collect and build their own dataset to evaluate the proposed methods. For fair and full comparison, we also list the evaluated scene/dataset for each method. It can be seen that the tightly coupling of VO and HD-MRL [8] improves the accuracy in long-term localization. The learning-based methods effectively solve the data association problem between 2D images and HD map elements, thus outperforming traditional methods, and they can localize without the requirements of online VO threads. BEV-Locator [234] estimates pose by regression whereas EgoVM [234] relies on exhaustively searching pose candidates, the latter solution achieves better performance. It should be noticed that HD map is a very compact format of map representation. In MLVHM [9], the HD maps stored in ASCII format only requires about 50 KB storage space per kilometer (km), compared to about 600 MB/km of the raw point cloud map. It is not the only case that the map size of EgoVM is 0.35 MB/km compared to 5.92 MB/km of DA4AD [10]. Therefore, the compactness and accuracy of HD map makes it a popular choice for localization in AD.
TABLE IX: Localization results on Cambridge Landmarks dataset [27].
@@
VLAD [29] 2.80,5.7° 4.01,7.1° 1.11,7.6° 2.31,8.0°
VLAD+Inter [306] 1.48,4.5° 2.68,4.6° 0.90,4.3° 1.62,6.1°
@@
EssNet [107] 0.76,1.9° 1.39,2.8° 0.84,4.3° 1.32,4.7°
NC-EssNet [107] 0.61,1.6° 0.95,2.7° 0.70,3.4° 1.12,3.6°
RelocGNN[256] 0.48,1.0° 1.14,2.5° 0.48,2.5° 1.52,3.2°
@@
PoseNet [27] 1.92,5.40° 2.31,5.38° 1.46,8.08° 2.65,8.48°
BayesianPN [241] 1.74,4.06° 2.57,5.14 ° 1.25,7.54° 2.11,8.38°
LSTM-PN [242] 0.99,3.65° 1.51,4.29° 1.18,7.44° 1.52,6.68°
MapNet [244] 1.07,1.89° 1.94,3.91° 1.49,4.22° 2.00,4.53°
@@
MSPN [249] 1.73,3.65° 2.55,4.05° 2.92,7.49° 2.67/6.18°
MS-Trans[307] 0.83,1.47° 1.81,2.39° 0.86,3.07° 1.62,3.99°
c2f-MS-Trans [251] 0.70,2.69° 1.48,2.94 ° 0.59,2.88° 1.14,3.88°
@@
DSAC [255] 0.30,0.5° 0.33,0.6° 0.09,0.4° 0.55,1.6°
DSAC++ [261] 0.18,0.3° 0.20,0.3° 0.06,0.3° 0.13,0.4°
$*{}^{*}$ [256] 0.18,0.3° 0.21,0.4° 0.05,0.3° 0.15,0.5°
SANet [43] 0.32/0.5° 0.32/0.5° 0.10/0.5° 0.16/0.6°
@@
DFNet [277] n.a. 0.46,0.87° 0.16,0.59° 0.50,1.49°
LENS [278] 0.33,0.5° 0.44,0.9° 0.27,1.6° 0.53,1.6°
@@
NeRF-Loc [275] 0.11,0.2° 0.18,0.4° 0.04,0.2° 0.07,0.2°
NN-MRL: Then, we discuss about MRL with learnt implicit map. According to the localization results on medium-scale Cambridge Landmarks dataset [27] in Tab. IX and small-scale 7Scenes dataset [27] in Tab. X, we find that the localization accuracy of APR methods is not pleased, and they performs much worse than VL-MRL methods. When enhanced by NeRF, the performances of APR methods can be significantly improved, and achieve comparable performances with VL-MRL methods. As a surprise, we find that SCR methods work well and achieve impressive localization accuracy in both medium-scale outdoor and small-scale indoor scenes. DSAC [255] and its extended works [261, 256] even work better than VL-MRL methods in some cases. SCR may be a promising trend in MRL research in the future. But its heavy cost of training should be considered. As for NeRF, it can significantly boost APR methods [277, 276, 278] and it also can work well as a pose estimator [275]. We are still waiting for the usage of NeRF in large-scale and long-term localization tasks since building NeRF for unbounded scene is a much more challenging problem compared to indoor scene.
Summary: Finally, we compare many MRL methods with various kind of scene map on outdoor Cambridge Landmarks [27] and indoor 7Scenes [27], as listed in Tab. IX and Tab. X. In outdoor scene, the rankings of MRL methods with regard to localization accuracy should be: NeRF pose estimator [275] $>$ I2P-MRL [213] $>$ SCR $>$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. And in indoor scene, the rankings of MRL methods are slightly changed as: NeRF pose estimator [275] $\approx$ I2P-MRL [213] $\approx$ SCR $\approx$ APR+NeRF $\approx$ VL-MRL $>$ RPR $>$ APR $>$ VPR. Please note that this general conclusion is drawn from the limited experimental results in this survey, there may be some isolated examples in MRL research.
It can be seen that in small-scale indoor scene, although APR and RPR work worse than other MRL methods, they can also achieve pleased localization performance. Given the low requirement of built map in APR and RPR, they can be used in some applications that do not need high-precision localization, such as VR. VPR only retrieve visually similar reference image and regard the reference pose as an approximated query pose, so its localization accuracy is very limited when reference images are few in outdoor scene. In indoor scene, most MRL methods with various kinds of scene map achieve comparable localization performance. But in outdoor scene, the gaps among different methods are more significant. Using NeRF as pose estimator like NeRF-Loc [275], I2P-MRL [213], SCR, and APR+NeRF seem to be promising ways in MRL researches. Adapting NeRF-based localization to large-scale unbounded scenes, reducing the requirement of high-precision point cloud map for I2P-MRL, reducing the training cost of SCR should be explored to make the MRL methods usable.
IX Opening Discussion
IX-A Are Explicit 3D Models Necessary in Visual Localization?
Traditional MRL methods require an explicit 3D scene map, e.g., visual landmarks map in VL-MRL, point cloud map in I2P-MRL, and vectorized map in HD-MRL. An explicit 3D model of the scene appears to be necessary in the early stage of MRL research. However, with the recent advancements in MRL, some techniques seem no longer require an explicit 3D model, such as VPR, RPR, and NN-MRL. Therefore, a question naturally arises: are explicit 3D models necessary for accurate visual localization?
As shown in Tab. IX and X, it can be seen that some 2D image-based methods without explicit map, like VPR and RPR, perform worse than 3D model-based method with explicit map, such as VL-MRL and I2P-MRL. However, 2D image-based methods with a learnt implicit map, i.e., NN-MRL, can achieve comparable or even better performances. Specially, NeRF-Loc [275] has been found to localize with centimeter-level accuracy, surpassing VL-MRL and I2P-MRL by a considerable margin. Thus, it is believed that the learnable implicit map has the potential to be a superior alternative to the explicit map for accurate localisation. The explicit 3D map is not strictly necessary in current DL age.
TABLE X: Localization results on 7Scenes dataset [27].